# Viewpoint Consistency in 3D Generation via Training-free Attention and CLIP Guidance

### v2

@@ 2.1 Text-to-3D Generation
Subsequent improvements to DreamFusion have focused on three areas: generation quality, speed, and the Janus problem [42, 21, 20, 4, 18, 12, 23, 19, 44, 5]. Efforts to improve quality mainly revolve around optimizing SDS, as seen in works like ProlificDreamer [42] and LucidDreamer [20], which analyze how SDS leads to mode collapse and excessive smoothing. Speed improvements typically involve optimizing the 3D representation and the generation framework. For example, Magic3D [21] introduces a two-stage optimization strategy and incorporates DMTet [35] for fast surface optimization, while DreamGaussian [41] employs 3D Gaussian Splatting [16] to accelerate generation. The Janus problem, which concerns generating consistent front and back views, is the primary focus of our research. Previous solutions rely on fine-tuning diffusion models to obtain multi-view capabilities [23, 36]. In contrast, our work takes a different approach by analyzing the inherent generative bias of diffusion models. We achieve superior results without compromising the model’s generalization ability.
@@ 2.2 Diffusion Model
where $f(x,t)$ is the drift coefficient, $g(t)$ is the diffusion coefficient, and $dw$ represents the differential of the Wiener process. In the reverse process, $\bar{\mathbf{w}}$ is the Wiener process when time flows backward from $T$ to 00 [39].
@@ 3.1 Rethink Janus Problem
where $p_{t}(x)$ is the time-evolved data distribution. This equation represents how the original data distribution $p_{0}(x)$ changes over time $t$ from 00 to $T$, eventually converging to a standard Gaussian distribution. For the reverse SDE, by letting $s=T-t$, we obtain a forward process from $s=0$ to $T$. Assuming the model’s predicted score is accurate, the Fokker-Planck equation for the reverse SDE becomes:
@@
which describes the evolution of the Gaussian distribution over time $s$ from 00 to $T$, corresponding to $t$ decreasing from $T$ to 00. This is essentially the reverse of Eq. 5. See supplementary for details.
@@ 3.3 CLIP Pruning
where $e_{\text{image}}$ and $e_{\text{text}}$ represent the embeddings of $\bar{x}_{0}$ and $P_{v}$, respectively. The dot product measures semantic alignment, and $\sigma$ is normalized to a range between $-1$ and $1$. If $\sigma$ falls below a specified threshold $\tau$, it indicates that the current generation provides incorrect guidance for the 3D model under the specified viewpoint In such cases, the update step is skipped to prevent the inclusion of misleading guidance that could cause the Janus problem.
@@ 5 Conclusion
Future Work. Having identified the Janus Problem as a generation bias caused by long-tailed distributions, future work could employ strategies that address long-tailed distribution issues to further alleviate the problem. ]

### v3

@@ 2.1 Text-to-3D Generation
Subsequent improvements to DreamFusion have focused on three areas: generation quality, speed, and the Janus problem [40, 19, 18, 4, 12, 21, 5]. Efforts to improve quality mainly revolve around optimizing SDS, as seen in works like ProlificDreamer [40] and LucidDreamer [18], which analyze how SDS leads to mode collapse and excessive smoothing. Speed improvements typically involve optimizing the 3D representation and the generation framework. For example, Magic3D [19] introduces a two-stage optimization strategy and incorporates DMTet [33] for fast surface optimization, while DreamGaussian [39] employs 3D Gaussian Splatting [16] to accelerate generation. The Janus problem, which concerns generating consistent front and back views, is the primary focus of our research. Previous solutions rely on fine-tuning diffusion models to obtain multi-view capabilities [21, 34]. In contrast, our work takes a different approach by analyzing the inherent generative bias of diffusion models. We achieve superior results without compromising the model’s generalization ability.
@@ 2.2 Diffusion Model
where $f(x,t)$ is the drift coefficient, $g(t)$ is the diffusion coefficient, and $dw$ represents the differential of the Wiener process. In the reverse process, $\bar{\mathbf{w}}$ is the Wiener process when time flows backward from $T$ to 0 [37].
@@ 3.1 Rethink Janus Problem
where $p_{t}(x)$ is the time-evolved data distribution. This equation represents how the original data distribution $p_{0}(x)$ changes over time $t$ from 0 to $T$, eventually converging to a standard Gaussian distribution. For the reverse SDE, by letting s=Tts=T-t, we obtain a forward process from $s=0$ to $T$. Assuming the model’s predicted score is accurate, the Fokker-Planck equation for the reverse SDE becomes:
@@
which describes the evolution of the Gaussian distribution over time $s$ from 0 to $T$, corresponding to $t$ decreasing from $T$ to 0. This is essentially the reverse of Eq. 5. See supplementary for details.
@@ 3.3 CLIP Pruning
where $e_{\text{image}}$ and $e_{\text{text}}$ represent the embeddings of $\bar{x}_{0}$ and $P_{v}$, respectively. The dot product measures semantic alignment, and $\sigma$ is normalized to a range between 1-1 and $1$. If $\sigma$ falls below a specified threshold $\tau$, it indicates that the current generation provides incorrect guidance for the 3D model under the specified viewpoint In such cases, the update step is skipped to prevent the inclusion of misleading guidance that could cause the Janus problem.
@@ 5 Conclusion
Future Work. Having identified the Janus Problem as a generation bias caused by long-tailed distributions, future work could employ strategies that address long-tailed distribution issues to further alleviate the problem.